# Lambert’s problem in orbital dynamics: a self–contained introduction

### v1

@@ Abstract
Lambert’s problem is a classical boundary value problem in analytical mechanics. It arises when trying to determine the energy required to place a particle, subject to a central gravitational potential, in a free fall trajectory connecting two given points on a desired travel time. Due to its mathematical beauty and its relevance in aerospace engineering, it has been and remains the object of attention of countless engineers, mathematicians (pure and applied), and physicists seeking to produce efficient solution algorithms. In this expository article, didactic in nature, we present a unified and comprehensive derivation that assumes only a minimal background in physics and mathematics. We focus on the simplest two–body case and carefully develop the argument for elliptical trajectories. The goal is to provide a single reference that can serve as an accelerated introduction for students and researchers interested in a quick introduction to the subject.
Keywords:
Classical mechanics Orbital transfer Orbital dynamics Keplerian dynamics Motion under a central potential Lambert problem.
MSC:
70-01 85-01 70M20 70F16 70F05.
@@ 1 Introduction
The goal of this expository article is didactic: we aim to provide a unified reference for readers who are comfortable with calculus, geometry and differential equations—but may not necessarily have strong background in physics—and are interested in the field of orbital dynamics in general and the Lambert problem in particular. All of the topics in this text are available in the literature—and have been available for a long while—however, a novice reader often has spend a significant amount of time and effort jumping from one reference to another just to piece together a reasonably strong background before being able to tackle research–level questions. This was in fact the experience the two first authors—graduate students in applied mathematics interested in the application of surrogate methods for the solution of Lambert’s problem Baloglou2025 ; BaGiSa2025 —and this text grew out of their background–building efforts. Given the instructional goal and our target readership, we provide significantly more details in the calculations than what a typical research article—and even some textbooks—would do. We believe that this approach (and this reference itself) will be especially useful to students—and even researchers—in physics, astronomy, aerospace engineering and applied mathematics who find themselves in need of an accelerated introduction to the subject.
@@
    The symbol :=assign:=:= denotes the fact that the quantity on the left is defined to be equal to the quantity on the right.
@@
    Differentiation with respect to time will often (but not always) be denoted by a dot on top of a variable; the number of dots denotes the order of the derivative, i.e. ${\displaystyle\dot{x}:=\tfrac{d}{dt}x}$ and ${\displaystyle\ddot{x}:=\tfrac{d^{2}}{dt^{2}}x}$.
@@ 2.1 Conic sections
Apollonian conics.
There are several equivalent definitions of the conic sections. Perhaps the most well–known are the association between them and planar intersections of cones, and the characterization in terms of quadratic forms. However, the most useful one for our purposes will be the definition due to the Greek geometer Apollonius of Perga Apollonius . The reason why we will use this definition is that it encompasses almost all of the definitions for conics in one single condition, in a way that is equivalent to the standard definition in terms of quadratic forms (GlStOd2016, ,  Theorem 2.1).
@@
is called a conic with associated focus $F$ and directrix $L$. Moreover, we say that the conic $\mathcal{C}$ is: an ellipse, if $10<e<$ 1, a parabola if $e=1$, or a hyperbola if $e>1$.
The line connecting $F$ with its directrix at a right angle (depicted in solid black in Figure 1) defines an axis of symmetry of the conic, known as the major or principal axis. The point $V$ along the segment $FL$ and belonging to the conic is known as the vertex. As we shall now see, ellipses and hyperbolas have a second axis of symmetry and therefore a second focus and directrix, however the eccentricity (with respect to the second pair focus/directrix) remains unchanged. The following argument is adapted from (GlStOd2016, ,  Lemma 2.1.1).
Let’s consider a directrix/focus pair where the directrix $L$ is oriented in the vertical direction and the focus $F$ is located to its left at some distance $|FL|$. Placing a Cartesian frame of reference with $x$ axis parallel to the major axis of symmetry and with its y axis parallel to the directrix $L$, we see that the focus has coordinates $(-|FL|,0)$. Therefore, for a point with coordinates $(x,y)$, the condition in Definition (1) can be expressed as
@@
which, in the elliptic case $10<e<$ 1, yields the value of the $y$ coordinate of the point in the ellipse located vertically above/below the center $O$. The semi–minor axis is defined as the distance
@@
Figure 1: Left: Conical sections for different values of the eccentricity $e$. Shades of red represent values of $10<e<$ 1 resulting in ellipses; the color fades towards one. Shades of blue represent values of $e>1$ resulting in hyperbolae with the color fading towards one. The parabola, plotted in violet, is the limiting case $e=1$. The directrix $L$ is plotted as a dashed line, the focus $F$ is marked by an open circle. Right: Relevant geometric markers for ellipses and hyperbolae.
The circle.
The reader must have noticed that the circle is missing from the definition above. We will now see that it can be understood as a special case of an ellipse with eccentricity $e=0$. However, if we simply let $e=0$ in Definition 1, we would only recover the degenerate case of a single point: the focus. Any attempt at using Equation (1) to obtain the circle as a limiting case of the Apollonian definition will face the same shortcoming. Hence, to include the circle as a limiting case, we will have to make the additional assumption that the length of the semi–major axis, $a$, is a fixed predetermined parameter independent from the eccentricity.
@@
Focal equation of a conic.
Placing a reference frame at the focus $F$ with the horizontal axis parallel to the line segment $FL$, it is possible to describe the position of a point $P$ in terms of its distance, $r$, from the focus and the angle, $\theta$, that the vector connecting $F$ to $P$ makes with the horizontal axis. For historical reasons, within the context of celestial mechanics, the angle $\theta$ is known as the true anomaly.
@@
Our derivation of equations (8) used the geometric construction from Figure 2, where the origin of the coordinate system is located at the right focus of the ellipse: the one labeled as $F_{1}$ in the top right panel of Figure 1, and the directrix is the line $x=|FL|$. However, it is possible and equally valid to place the frame of reference at the left focus (labeled $F_{2}$ in Figure 1) so that the directrix is the line $x=-|FL|$. In that case, if the angle is still measured from the positive horizontal axis, the equations derived by the same procedure will be analogous to (8) with the only difference being that the eccentricity in the denominator will appear with the opposite sign. Hence, depending on the references, the denominator of (8) can be equal to $1+e\cos\theta$, as in our case, or to $1-e\cos\theta$. Some references go as far as writing the denominator as $1\pm e\cos\theta$. The reader should be aware of this possible discrepancy and how to resolve it.
@@ 2.2 Elementary concepts from mechanics
The following elements from theoretical mechanics can be found in any standard reference, such as Morin2012 ; Scheck2010 ; ThMa2014 . For the benefit of the reader with no previous familiarity with physics, we collect here just concepts that will be absolutely essential for our exposition.
Governing laws.
We are ultimately interested in describing the motion of a point particle (i.e. one that can be idealized as occupying no volume) with mass $m$ under the action of a gravitational force $\bm{F}$ produced by an attracting body of mass $M$. We will refer to the second body as the attractive center and will also consider it to be a point. We will place a reference frame at the location of the attracting center and will determine the position of the particle of mass $m$ by the vector $\bm{r}$ anchored at the origin connecting the two particles. We will denote the magnitude of the position vector by $r:=|\bm{r}|$. As is customary in mechanics and dynamical systems, we will use a dot to represent time differentiation, and will denote the velocity of the particle by $\dot{\bm{r}}$. Under these circumstances, the mathematical description of the motion of the particle will be determined by
[…]
@@
A priori, the expression above is a system of three second order differential equations (one for each component of the position vector $\bm{r}$) that, given the appropriate initial conditions, completely determines the movement of a particle under the influence of the attractive center. Rather than setting out to integrate these equations, in what follows we will exploit the physical properties of the system to reduce the system first into a two dimensional second order vector problem, and then into a first order scalar problem.
Conservation of angular momentum.
For a point particle with mass $m$, the vector quantity
@@
is known as the angular momentum. The symbol “$\times$” above denotes the standard vector or cross product satisfying for all $\bm{u},\bm{v}\in\mathbb{R}^{3}$:
@@
$\dot{\bm{L}}=\frac{d}{dt}\left(\bm{r}\times m\dot{\bm{r}}\right)=\dot{\bm{r}}\times m\dot{\bm{r}}+\bm{r}\times m\ddot{\bm{r}}=\dot{\bm{r}}\times m\dot{\bm{r}}+\bm{r}\times\left(-\frac{GMm}{r^{3}}\bm{r}\right)=\bm{0},$
where we used (10) and the fact that the cross product vanishes when its arguments are multiples of each other. The result above is known as the conservation of angular momentum and has long reaching consequences. The first one is that, due to $\bm{L}$ being normal to the plane of motion, {adjustwidth}3cm0cm under the action of a gravitational potential, the motion of a particle is constrained to the fixed plane determined by its initial position and velocity. This fact has the effect of reducing the dimensionality of the problem from three to two dimensions, allowing us to use a polar coordinate system.
Position, velocity, and acceleration in polar coordinates
In view of the first geometric consequence of the conservation of angular momentum, we will assume that the plane of motion coincides with the Euclidean $xy$ plane and will express the position vector $\bm{r}$ in polar coordinates in terms of its magnitude $r:=|\bm{r}|$ and the angle $\theta$ that it makes with respect to the positive $x$–axis.
@@
$\frac{d}{dt}\widehat{\bm{r}}=\dot{\theta}\widehat{\bm{\theta}}\qquad\text{ and}\qquad\frac{d}{dt}\widehat{\bm{\theta}}=-\dot{\theta}\widehat{\bm{r}}.$ (12)
@@
Figure 3: Left: For a fixed angle $\theta$ the polar basis vectors are orthogonal and point in the direction of growth $(1+\Delta)\bm{r}$ (in gray) and $\theta+\Delta\theta$ (in red). Right: The velocity of a particle can be decomposed in radial and angular components.
Conservation of energy.
The kinetic energy of a particle with mass $m$ moving with velocity $\dot{\bm{r}}$ is defined as
@@
where we used Equation (13b) and the fact that the vectors $\widehat{\bm{r}}$ and $\widehat{\bm{\theta}}$ are perpendicular. The first term above is due to the radial component of the velocity—and is therefore known as the radial kinetic energy, while the second one is due to the angular component—and is referred to as the rotational energy.
@@
=\,= $\frac{1}{2}m(2\dot{r}\ddot{r}+2r\dot{r}\dot{\theta}^{2}+2r^{2}\dot{\theta}\ddot{\theta})+\frac{GMm}{r^{2}}\dot{r}$
=\,= $m(\dot{r}\ddot{r}+r\dot{r}\dot{\theta}^{2}+r^{2}\dot{\theta}\ddot{\theta})+\frac{GMm}{r^{2}}\dot{r}.$ (14)
@@ 3.1 Equations of motion
is constant. These two equations involve only first derivatives of the radial and angular coordinates and for that reason are sometimes referred to as first integrals of motion. Solving for $\dot{r}$ and $\dot{\theta}$ in the expressions above yields the first order, nonlinear system of differential equations:
@@ 3.2 Gravitational orbits
The exception to this, however, is the case when $L=0$, as in that case (17a) is independent of $r$—note that equation (17b) depends exclusively on $r$. We will study this case first. If the angular velocity $\dot{\theta}$ vanishes at any given time, then the angular momentum would vanish as well. Therefore, by the conservation property encoded in equation (17a), this would imply that $\dot{\theta}=0$ for all times and the particle would describe a straight line at a constant angle $\theta_{0}=\theta(0)$.
On the other hand, if the angular speed $\dot{\theta}$ is not zero initially, then the conservation of angular momentum implies that it will never vanish—as the left hand side of (17a) would then be a positive constant. In that case $r$ and $\theta$ are functionally dependent and we can assume an implicit relationship of the form $r=r(\theta(t))$. Using the chain rule we obtain:
@@
$-\frac{du}{d\theta}=\sqrt{\frac{2m}{L^{2}}\left(E+GMmu\right)-u^{2}}.$ (18)
@@
and realizing that ${\displaystyle\frac{du}{d\theta}=\frac{dy}{d\theta}}$, the equation (18) results in
@@
Equation (19a) describes all the possible trajectories for particles with non–zero angular momentum moving under a gravitational potential [^1] We discussed the case of zero angular momentum at the beginning of this section, which results in straight lines connecting the initial position of the particle with the attractive center.. The attentive reader will note that (19a) is exactly of the form (8) that we arrived at when studying conic sections. This allows us to state the following remarkable fact: {adjustwidth}1.5cm0cm The only possible trajectories for a particle under a central gravitational potential are conic sections. From a purely geometric point of view, the particular conic is determined by the eccentricity. Equation (19b) connects the geometry of the problem to the physical parameters by quantifying the way in which the eccentricity $e$ depends on the particles energy $E$ and angular momentum $L$. Concretely, expressing $p$ in terms of the angular momentum and equating (19a) with (8b) we can connect the geometric parameters $e$ and $a$ with the physical parameters, resulting in
@@
    $\boxed{\bm{L=0\,.}\text{{ Straight lines.}}}$. If the angular momentum is zero, then (17a) implies that the angle remains constant and the particle’s orbit is a straight line.
@@
      The expressions for the radius of circular orbits and the semi–major axis of elliptic orbits that we just derived, confirm that a particle with no angular momentum necessarily collapses into the attractive center. They also reveal that, for a fixed initial angular momentum, more massive the bodies (i.e. those for which the product $Mm^{2}$ is larger), are bound to remain closer to the attractive center.
@@
      $\lim_{r\to\infty}\dot{\theta}=0\qquad\text{ and }\qquad\lim_{r\to\infty}\dot{r}=\dot{r}_{\infty}<\infty.$
@@
        $\boxed{e=1\,.\,\text{{Parabolic orbits.}}}$. If the potential term equals the kinetic term, the total energy will be $E=0$. Equation (19b) implies that $e=1$ and the particle follows a parabolic orbit. Comparing (19a) and (8a) it follows that the distance between the directrix of the parabola and the attractive center is
@@
        while, as the angle $\theta$ approaches $\pi$ or $-\pi$, equation (19a) implies that $r\to\infty.$ We note that as $r\to\infty$ the potential energy vanishes. The limit condition (22) and the fact that $E=0$ imply that the residual speed at infinity for a particle on a parabolic trajectory is $\dot{r}_{\infty}=0$.
@@
        $\boxed{e>1\,.\,\text{{Hyperbolic orbits.}}}$. Finally, when the kinetic energy exceeds the potential energy the eccentricity of the trajectory will be greater than one, and the particle will follow a hyperbolic arc. In this case the minimum distance to the attractive center will be given by
@@
In the elliptic and hyperbolic cases, the numerators from (8b) and (19a) can be equated and the form of $e$ in (19b) can be used to obtain the expression for the semi–major axis $a$ in terms of the physical parameters

### v2

@@ Abstract
Lambert’s problem is a classical boundary value problem in analytical mechanics. It arises when trying to determine the energy required to place a particle, subject to a central gravitational potential, in a free fall trajectory connecting two given points on a desired travel time. Due to its mathematical beauty and its relevance in aerospace engineering, it has been and remains the object of attention of countless engineers, mathematicians (pure and applied), and physicists seeking to produce efficient solution algorithms. In this expository article, didactic in nature, we present a unified and comprehensive derivation that assumes only a minimal background in physics and mathematics. We focus on the simplest unperturbed case and carefully develop the argument for elliptical trajectories. The goal is to provide a single reference that can serve as an accelerated introduction for students and researchers interested in a quick introduction to the subject.
@@ 1 Introduction
The goal of this expository article is didactic: we aim to provide a unified reference for readers who are comfortable with calculus, geometry and differential equations—but may not necessarily have strong background in physics—and are interested in the field of orbital dynamics in general and the Lambert problem in particular. All of the topics in this text are available in the literature—and have been available for a long while—however, a novice reader often has to spend a significant amount of time and effort jumping from one reference to another just to piece together a reasonably strong background before being able to tackle research–level questions. This was in fact the experience the two first authors—graduate students in applied mathematics interested in the use of surrogate methods for the solution of Lambert’s problem Baloglou2025 ; BaGiSa2025 ; this text grew out of their background–building efforts. Given the instructional goal and our target readership, we provide significantly more details in the calculations than what a typical research article—and even some textbooks—would do. We believe that this approach (and this reference itself) will be especially useful to students—and even researchers—in physics, astronomy, aerospace engineering and applied mathematics who find themselves in need of an accelerated introduction to the subject.
@@
    The symbol :=:= denotes the fact that the quantity on the left is defined to be equal to the quantity on the right.
@@
    Differentiation with respect to time will often (but not always) be denoted by a dot on top of a variable; the number of dots denotes the order of the derivative, i.e. $\dot{x}:=\tfrac{d}{dt}x$} and $\ddot{x}:=\tfrac{d^{2}}{dt^{2}}x$}.
@@ 2.1 Conic sections
Apollonian conics. There are several equivalent definitions of the conic sections. Perhaps the most well–known are the association between them and planar intersections of cones, and the characterization in terms of quadratic forms. However, the most useful one for our purposes will be the definition due to the Greek geometer Apollonius of Perga Apollonius . The reason why we will use this definition is that it encompasses almost all of the definitions for conics in one single condition, in a way that is equivalent to the standard definition in terms of quadratic forms (GlStOd2016, ,  Theorem 2.1).
@@
is called a conic with associated focus $F$ and directrix $L$. Moreover, we say that the conic $\mathcal{C}$ is: an ellipse, if $0<e<1$, a parabola if $e=1$, or a hyperbola if $e>1$.
The line connecting $F$ with its directrix at a right angle (depicted in solid black in Figure 1) defines an axis of symmetry of the conic, known as the major or principal axis. The point $V$ along the segment $FL$ and belonging to the conic is known as the vertex. As we shall now see, ellipses and hyperbolas have a second axis of symmetry and therefore a second focus and directrix, however the eccentricity (with respect to the second pair focus/directrix) remains unchanged. The following argument has been adapted from (GlStOd2016, ,  Lemma 2.1.1).
Let’s consider a directrix/focus pair where the directrix $L$ is oriented in the vertical direction and the focus $F$ is located to its left at some distance $|FL|$. Placing a Cartesian frame of reference with $x$ axis parallel to the major axis of symmetry and with its $y$ axis parallel to the directrix $L$, we see that the focus has coordinates (|FL|,0)(-|FL|,0). Therefore, for a point with coordinates $(x,y)$, the condition in Definition (1) can be expressed as
@@
which, in the elliptic case $0<e<1$, yields the value of the $y$ coordinate of the point in the ellipse located vertically above/below the center $O$. The length of the semi–minor axis is defined as the distance
@@
Figure 1: Left: Conical sections for different values of the eccentricity $e$. Shades of red represent values of $0<e<1$ resulting in ellipses; the color fades towards one. Shades of blue represent values of $e>1$ resulting in hyperbolae with the color fading towards one. The parabola, plotted in violet, is the limiting case $e=1$. The directrix $L$ is plotted as a dashed line, the focus $F$ is marked by an open circle. Right: Relevant geometric markers for ellipses and hyperbolae.
The circle. The reader must have noticed that the circle is missing from the definition above. We will now see that it can be understood as a special case of an ellipse with eccentricity $e=0$. However, if we simply let $e=0$ in Definition 1, we would only recover the degenerate case of a single point: the focus. Any attempt at using Equation (1) to obtain the circle as a limiting case of the Apollonian definition will face the same shortcoming. Hence, to include the circle as a limiting case, we will have to make the additional assumption that the length of the semi–major axis, $a$, is a fixed predetermined parameter independent from the eccentricity.
@@
Focal equation of a conic. Placing a reference frame at the focus $F$ with the horizontal axis parallel to the line segment $FL$, it is possible to describe the position of a point $P$ in terms of its distance, $r$, from the focus and the angle, $\theta$, that the vector connecting $F$ to $P$ makes with the horizontal axis. For historical reasons, within the context of celestial mechanics, the angle $\theta$ is known as the true anomaly.
@@
Our derivation of equations (8) used the geometric construction from Figure 2, where the origin of the coordinate system is located at the right focus of the ellipse: the one labeled as $F_{1}$ in the top right panel of Figure 1, and the directrix is the line $x=|FL|$. However, it is possible and equally valid to place the frame of reference at the left focus (labeled $F_{2}$ in Figure 1) so that the directrix is the line x=|FL|x=-|FL|. In that case, if the angle is still measured from the positive horizontal axis, the equations derived by the same procedure will be analogous to (8) with the only difference being that the eccentricity in the denominator will appear with the opposite sign. Hence, depending on the references, the denominator of (8) can be equal to $1+e\cos\theta$, as in our case, or to $1-e\cos\theta$. Some references go as far as writing the denominator as $1\pm e\cos\theta$. The reader should be aware of this possible discrepancy and how to resolve it.
@@ 2.2 Elementary concepts from mechanics
The following elements from theoretical mechanics can be found in any standard reference, such as Morin2012 ; Scheck2010 ; ThMa2014 . For the benefit of the reader with no previous familiarity with physics, we collect here just the concepts that will be absolutely essential for our exposition.
Governing laws. We are ultimately interested in describing the motion of a point particle (i.e. one that can be idealized as occupying no volume) with mass $m$ under the action of a gravitational force $\bm{F}$ produced by an attracting body of mass $M$. We will refer to the body with mass $M$ as the attractive center and will also consider it to be a point. We will place a reference frame at the location of the attracting center and will determine the position of the particle of mass $m$ by the vector $\bm{r}$, anchored at the origin, connecting the two particles. We will denote the magnitude of the position vector by $r:=|\bm{r}|$. As is customary in mechanics and dynamical systems, we will use a dot to represent time differentiation, and will denote the velocity of the particle by $\dot{\bm{r}}$. Under these circumstances, the mathematical description of the motion of the particle will be determined by
$m\ddot{\bm{r}}=\,$ $\bm{F}$ $\text{(Newton's second law of motion)},$
$\bm{F}=\,$ $-\frac{GMm}{r^{3}}\bm{r}\qquad\qquad$ $\text{(Newton's universal law of gravitation)}.$
@@
A priori, the expression above is a system of three second order differential equations (one for each component of the position vector $\bm{r}$) that, given the appropriate initial conditions, completely determines the movement of a particle under the influence of the attractive center. Rather than setting out to integrate these equations, in what follows we will exploit the physical properties of the system to reduce it first into a two dimensional second order vector problem, and then into a first order scalar problem.
Conservation of angular momentum. For a point particle with mass $m$, the vector quantity
@@
is known as the angular momentum. The symbol $\times$” above denotes the standard vector or cross product satisfying for all $\bm{u},\bm{v}\in\mathbb{R}^{3}$:
@@
[…]
where we used the fact that the cross product vanishes when its arguments are multiples of each other. The result above is known as the conservation of angular momentum and has long reaching consequences. The first one is that, due to $\bm{L}$ being normal to the plane of motion, under the action of a gravitational potential, the motion of a particle is constrained to the fixed plane determined by its initial position and velocity. This fact has the effect of reducing the dimensionality of the problem from three to two dimensions, allowing us to use a polar coordinate system.
Position, velocity, and acceleration in polar coordinates In view of the first geometric consequence of the conservation of angular momentum, we will assume that the plane of motion coincides with the Euclidean $xy$ plane and will express the position vector $\bm{r}$ in polar coordinates in terms of its magnitude $r:=|\bm{r}|$ and the angle $\theta$ that it makes with respect to the positive xx–axis.
@@
[…]
@@
Figure 3: Left: For a fixed angle $\theta$ the polar basis vectors are orthogonal and point in the direction of growth $(1+\Delta)\bm{r}$ (in gray) and $\theta+\Delta\theta$ (in red). Right: The velocity of a particle, $\dot{\bm{r}}$, can be decomposed in radial and angular components.
Conservation of energy. The kinetic energy of a particle with mass $m$ moving with velocity $\dot{\bm{r}}$ is defined as
@@
where we used Equation (13b) and the fact that the vectors $\widehat{\bm{r}}$ and $\widehat{\bm{\theta}}$ are perpendicular. The first term above is due to the radial component of the velocity—and is therefore known as the radial kinetic energy—while the second one is due to the angular component—and is referred to as the rotational energy.
@@
$=\,$ $\frac{1}{2}m(2\dot{r}\ddot{r}+2r\dot{r}\dot{\theta}^{2}+2r^{2}\dot{\theta}\ddot{\theta})+\frac{GMm}{r^{2}}\dot{r}$
$=\,$ $m(\dot{r}\ddot{r}+r\dot{r}\dot{\theta}^{2}+r^{2}\dot{\theta}\ddot{\theta})+\frac{GMm}{r^{2}}\dot{r}.$ (14)
@@ 3.1 Equations of motion
is constant. These two equations involve only first derivatives of the radial and angular coordinates and for that reason are sometimes referred to as first integrals of motion. Solving for $\dot{r}$ and $\dot{\theta}$ in the expressions above yields the first order nonlinear system of differential equations:
@@ 3.2 Gravitational orbits
The exception to this, however, is the case when $L=0$, as in that case (17a) is independent of rr—note that equation (17b) depends exclusively on $r$. We will study this case first. If the angular velocity $\dot{\theta}$ vanishes at any given time, then the angular momentum would vanish as well. Therefore, by the conservation property encoded in equation (17a), this would imply that $\dot{\theta}=0$ for all times and the particle would describe a straight line at a constant angle $\theta_{0}=\theta(0)$.
On the other hand, if the angular speed $\dot{\theta}$ is not zero initially, then the conservation of angular momentum implies that it will never vanish—as the left hand side of (17a) would then be a positive constant. In that case $r$ and $\theta$ are functionally dependent on each other and we can assume an implicit relationship of the form $r=r(\theta(t))$. Using the chain rule we obtain:
@@
$\frac{du}{d\theta}=\sqrt{\frac{2m}{L^{2}}\left(E+GMmu\right)-u^{2}}.$ (18)
@@
and realizing that $\frac{du}{d\theta}=\frac{dy}{d\theta}$}, the equation (18) results in
@@
Equation (19a) describes all the possible trajectories for particles with non–zero angular momentum moving under a gravitational potential [^1] We discussed the case of zero angular momentum at the beginning of this section, which results in straight lines connecting the initial position of the particle with the attractive center.. The attentive reader will note that (19a) is exactly of the form (8) that we arrived at when studying conic sections. This allows us to state the following remarkable fact: The only possible trajectories for a particle under a central gravitational potential are conic sections. From a purely geometric point of view, the particular conic is determined by the eccentricity. Equation (19b) connects the geometry of the problem to the physical parameters by quantifying the way in which the eccentricity $e$ depends on the particle’s energy $E$ and angular momentum $L$. Concretely, expressing $p$ in terms of the angular momentum and equating (19a) with (8b), we can connect the geometric parameters $e$ and $a$ with the physical parameters, resulting in
@@
    $\boxed{\bm{L=0\,.}\text{{ Straight lines.}}}$ If the angular momentum is zero, then (17a) implies that the angle remains constant and the particle’s orbit is a straight line.
@@
      The expressions for the radius of circular orbits and the semi–major axis of elliptic orbits that we just derived, confirm that a particle with no angular momentum necessarily collapses into the attractive center. They also reveal that, for a fixed initial angular momentum, more massive bodies (i.e. those for which the product $Mm^{2}$ is larger) are bound to remain closer to the attractive center.
@@
[…]
@@
        $\boxed{e=1\,.\,\text{{Parabolic orbits.}}}$ If the potential term equals the kinetic term, the total energy will be $E=0$. Equation (19b) implies that $e=1$ and the particle follows a parabolic orbit. Comparing (19a) and (8a) it follows that the distance between the directrix of the parabola and the attractive center is
@@
        while, as the angle $\theta$ approaches $\pi$ or $\pi$, equation (19a) implies that $r\to\infty.$ We note that as $r\to\infty$ the potential energy vanishes. The limit condition (22) and the fact that $E=0$ imply that the residual speed at infinity for a particle on a parabolic trajectory is $\dot{r}_{\infty}=0$.
@@
        $\boxed{e>1\,.\,\text{{Hyperbolic orbits.}}}$ Finally, when the kinetic energy exceeds the potential energy the eccentricity of the trajectory will be greater than one, and the particle will follow a hyperbolic arc. In this case the minimum distance to the attractive center will be given by
@@
In the elliptic and hyperbolic cases, the numerators from (8b) and (19a) can be equated, and the form of $e$ in (19b) can be used to obtain the expression for the semi–major axis $a$ in terms of the physical parameters